5. Two polaroids are oriented at the an to each other and a plane-polarized light is gle incident upon them. If the fraction n of the incident light gets through both of them, what was the angle 0 between the polarization direction of the light and the polarization axis of the first polaroid? After obtaining the formula for 0, calculate its numerical value for -3So
Answer #1

Let a plan polarized light of intensity I0 is incident on first polarizer.

Then from law of malus-

                I=I0 cos2(theta)

At second polarizer applying law of malus-

                I’=Icos2(phi)= I0 cos2(theta) cos2(phi)

Also n=I’/I0

                n= cos2(theta) cos2(phi)

                cos2(theta) = n/ cos2(phi)    

                theta=cos-1(n1/2/cos(phi) )

if phi=350 and n=0.1

then

                cos(350)=0.819

                theta=cos-1(n1/2/cos(phi) )

                theta= 67.29